# Rethinking MUSHRA: Addressing Modern Challenges in Text-to-Speech Evaluation

### v2

@@ 5 Rethinking MUSHRA
We summarize two issues identified in Section 4. First, reference-matching bias that arises when listeners rate systems that perform at or above the level of the reference lower than deserved due to their efforts to align system outputs with the reference during evaluation. Second, judgement ambiguity that arises when listeners rate a system on a single scale using broadly defined metrics like “naturalness”, leaving room for subjective interpretation of sub-criteria such as “prosody”, “voice quality”, “liveliness”, etc. leading to high variability in ratings. In response to this, we propose two refined variants of the MUSHRA test to address the identified challenges, as described below.
@@
MUSHRA-DG. The second variant, MUSHRA-DG (MUSHRA With Detailed Guidelines), introduces comprehensive guidelines to reduce the ambiguity in rating samples for naturalness. In this test, we present raters with scoresheets and a formula to arrive at MUSHRA scores systematically. Each rater was asked to mark the number of (i) mild pronunciation mistakes, (ii) severe pronunciation mistakes, (iii) unnatural pauses, speedups, or slowdowns, (iv) digital artifacts, (v) sudden energy fluctuations, and (vi) word skips. Further, raters were also asked to rate more perceptual measures such as (i) liveliness, (ii) voice quality, and (iii) rhythm on a continuous scale from 0-100. The detailed guidelines provided to raters to assess across each of these dimensions can be found in Appendix A.4. We analytically derive a MUSHRA naturalness score to raters using an intuitive formula with weights (provided in Appendix A.4) for different dimensions listed above. These weights can be tweaked depending on the specific use-case. For example, in a TTS application designed for audiobooks, where fluidity and expressiveness are crucial for user engagement, we might assign higher weights to liveliness and rhythm.
We understand that devising a scoring formula involves some subjectivity. To address this, we encouraged raters to review their evaluations and adjust their fine-grained scores if they felt the overall scores from the formula did not accurately reflect the differences they perceived between system pairs. Notably, we tracked these revisions and found that they occurred in only 1.3% of cases. This way, the final ratings better represented the raters’ true opinions about each system’s quality while also reducing any shortcomings that could have stemmed from the formula. More interestingly, we notice that the scores derived from the MUSHRA-DG test preserve the rankings obtained from the gold-standard Comparative Mean Opinion Score (CMOS) tests (Table 3), thus reinforcing the validity of our evaluations. Additionally, the variance in MUSHRA scores calculated using the formula is significantly lower, indicating reduced ambiguity in ratings and providing a clearer distinction between different systems’ performances.
@@ 6.2 Evaluations using MUSHRA-DG
While system rankings in closely matched scenarios can be debated, we argue that such cases highlight the need to focus on fine-grained differences. Understanding the specific contexts in which one model outperforms another provides valuable insights into system behavior and guides targeted improvements. Keeping this in mind, we subsequently discuss the prowess of the MUSHRA-DG test in fault isolation.
@@ 7 Conclusion
Our comprehensive study reveals significant shortcomings in the current use of the MUSHRA test for evaluating modern high-quality TTS systems. Through an extensive analysis involving 246,000 human ratings, we identified two primary issues: reference-matching bias and judgment ambiguity. To address these issues, we propose two refined variants of the MUSHRA test: MUSHRA-NMR, which omits explicit identification of the human reference, and MUSHRA-DG, which uses detailed guidelines to calculate MUSHRA scores systematically. Our findings indicate that both variants lead to more reliable evaluations, with MUSHRA-DG offering the additional benefit of fine-grained fault isolation during assessment. Through this work, we also release MANGO, a large human rating dataset, to further support research in this area.
@@ A.1 Instructions For MUSHRA
A.2 Sensitivity of MUSHRA in Tamil
@@
A.3 Visualizing MUSHRA Distributions
In Section 4, we discussed the distribution of MUSHRA scores across raters for Hindi using Figure 1. Similarly, in Figure 10 , we visualize the MUSHRA scores per rater across the three systems for Tamil. The Figure 11 and Figure 12, visualizes the MUSHRA scores for each utterance, averaged across raters, for Hindi and Tamil respectively.
A.4 Detailed Guidelines for MUSHRA-DG
@@
A.5 Fault Isolation in Tamil
Similar to Figure 5, we visualize the granular ratings across the objective and perceptual dimensions for Tamil in Figure 9.
@@ A.5 Fault Isolation in Tamil
A.6 Limitations
@@
Our study focuses on human evaluations for Hindi and Tamil, representing a major Indo-Aryan and Dravidian language, respectively. However, we did not extend our analysis to English, a widely spoken and diverse language. This limitation is due to the scope of our current research and resource constraints. We also wish to state that while the demographics table summarizes the age and gender of participants, details for 21 out of the 492 participants are missing due to system errors; this constitutes a minor fraction and we believe does not significantly affect the analysis. While this work focuses on identifying two key issues — reference-matching bias and judgment ambiguity — several other biases may also exist in TTS evaluations, which warrant further investigation. Future studies should include evaluations in English to generalize our findings across different language families and understand how language-specific characteristics might influence TTS evaluation outcomes. This broader analysis could provide more comprehensive insights into the applicability and robustness of our proposed MUSHRA variants across diverse linguistic contexts.
A.7 Ethical Considerations
@@
We prioritized ethical conduct throughout our research. The 492 human listeners involved in the study provided informed consent before participating in the evaluation, recruited through professional data annotation agencies. These agencies verified participant language proficiency for task relevance. We established an education criterion of completing grade 12 (Indian system) to ensure participants’ ability to accurately annotate audio content. Participants were compensated fairly for their time and expertise, following industry standards. They were also fully informed about the study nature, procedures, and their right to withdraw at any point without consequence. We ensured that our study adhered to ethical guidelines by obtaining approval from the Institutional Ethics Committee, which reviewed our methodology and confirmed compliance with ethical standards.
@@
We strived for inclusivity and bias mitigation. Participants came from diverse demographic backgrounds, and for Hindi and Tamil evaluations, we recruited only native speakers to capture the subtle linguistic and cultural nuances of each language. To minimize rater burden and bias in the new MUSHRA test variations, we prioritized user-friendliness and transparency in the design, providing clear guidelines.
@@
We release the evaluations dataset, which includes 246,000 human ratings, under CC-BY-4.0 license after careful consideration of privacy and ethical use. Identifiable information about the participants was anonymized to protect their privacy. We encourage the use of this dataset for advancing TTS evaluation metrics, emphasizing that it should be used responsibly and ethically, adhering to principles of transparency and fairness. Finally, we acknowledge that our study focuses on Hindi and Tamil, and we recognize the importance of extending such evaluations to other languages, including English, to generalize our findings. Future research should continue to explore these ethical dimensions, ensuring that the development and evaluation of TTS systems are conducted with respect for the diversity and rights of all participants involved.

### v3

@@ 4.6 Does adding more systems affect scores?
Table 5: Objective evaluation of TTS systems in Hindi and Tamil, measuring intelligibility, speech quality, and signal distortion.
System Hindi Tamil
STOI PESQ SI-SDR STOI PESQ SI-SDR
ANC 0.964 3.39 27.53 0.995 3.65 22.29
ST2 0.997 3.72 24.72 0.998 3.99 28.63
REF 0.997 3.93 27.76 0.998 4.08 30.25
VITS 0.998 4.11 30.17 0.997 3.98 25.63
FS2 0.999 4.17 28.89 0.998 4.09 26.74
4.7 Do objective metrics align with human perception?
To complement our subjective evaluations, we compute three widely used objective metrics — (i) PESQ (Perceptual Evaluation of Speech Quality), (ii) STOI (Short-Time Objective Intelligibility), and (iii) SI-SDR (Scale-Invariant Signal-to-Distortion Ratio), using TorchAudio-Squim (Kumar et al., 2023a). PESQ estimates speech quality by assessing distortions caused by noise and compression, while STOI measures intelligibility by evaluating how well speech information is preserved. SI-SDR, commonly used in speech enhancement, quantifies the level of distortion in a signal relative to a clean reference. These metrics offer quantifiable insights, but are primarily designed and validated on English datasets, making their applicability to languages like Hindi and Tamil uncertain.
Table 5 reveals weak alignment between MUSHRA scores and both PESQ and STOI, suggesting that these metrics do not fully capture perceptual speech quality. This misalignment may stem from their focus on intelligibility and signal distortions rather than higher-level perceptual attributes such as naturalness, prosody, and timbre. In contrast, SI-SDR exhibits a strong Pearson correlation of 0.76 with MUSHRA scores, indicating that the degree of signal distortion plays a key role in perceptual judgments, or that SI-SDR effectively captures aspects of TTS quality that influence human ratings. These findings highlight the limitations of existing objective metrics and the need for more refined evaluation methods that better align with perceptual judgments, either as complementary measures or as reliable alternatives to subjective ratings.
@@ 5 Rethinking MUSHRA
We summarize two issues identified in Section 4. First, reference-matching bias that arises when listeners rate systems that perform at or above the level of the reference lower than deserved due to their efforts to align system outputs with the reference during evaluation. Second, judgement ambiguity that arises when listeners rate a system on a single scale using broadly defined metrics like “naturalness”, leaving room for subjective interpretation of sub-criteria such as “prosody”, “voice quality”, “liveliness”, etc. leading to high variability in ratings. Findings from our post-evaluation survey (Appendix A.5) further confirm the presence of these challenges, highlighting the influence of reference-matching and the inconsistencies in how raters interpret evaluation criteria. In response to this, we propose two refined variants of the MUSHRA test to address the identified challenges, as described below.
@@
MUSHRA-DG. The second variant, MUSHRA-DG (MUSHRA With Detailed Guidelines), introduces comprehensive guidelines to reduce the ambiguity in rating samples for naturalness. In this test, we present raters with scoresheets and a formula to arrive at MUSHRA scores systematically. Each rater was asked to mark the number of (i) mild pronunciation mistakes, (ii) severe pronunciation mistakes, (iii) unnatural pauses, speedups, or slowdowns, (iv) digital artifacts, (v) sudden energy fluctuations, and (vi) word skips. Further, raters were also asked to rate more perceptual measures such as (i) liveliness, (ii) voice quality, and (iii) rhythm on a continuous scale from 0-100. The detailed guidelines provided to raters to assess across each of these dimensions can be found in Appendix A.6. We analytically derive a MUSHRA naturalness score for raters using an intuitive formula with weights (provided in Appendix A.6) for different dimensions listed above. These weights can be tweaked depending on the specific use-case. For example, in a TTS application designed for audiobooks, where fluidity and expressiveness are crucial for user engagement, we might assign higher weights to liveliness and rhythm.
We understand that devising a scoring formula involves some subjectivity. To address this, we encouraged raters to review their evaluations and adjust their fine-grained scores if they felt the overall scores from the formula did not accurately reflect the differences they perceived between system pairs. Notably, we tracked these revisions and found that they occurred in only 1.3% of cases. This suggests that raters likely found the computed scores to align with their assessments. Such alignment is reasonable, as the test design and the formula, including its specific weights and penalties, were refined through iterative evaluations and sustained engagement with raters over several months. Throughout this process, adjustments to the guidelines were made based on direct feedback and practical insights from raters, while also being guided by the intuition of TTS practitioners.
More interestingly, we notice that the scores derived from the MUSHRA-DG test preserve the rankings obtained from the gold-standard Comparative Mean Opinion Score (CMOS) tests (Table 3), thus reinforcing the validity of our evaluations. Additionally, the variance in MUSHRA scores calculated using the formula is significantly lower, indicating reduced ambiguity in ratings and providing a clearer distinction between different systems’ performances.
@@ 6.2 Evaluations using MUSHRA-DG
While system rankings in closely matched scenarios can be debated, we argue that such cases highlight the need to focus on fine-grained differences. Understanding the specific contexts in which one model outperforms another provides valuable insights into system behavior and guides targeted improvements. Keeping this in mind, we subsequently discuss the effectiveness of the MUSHRA-DG test in fault isolation.
@@
Figure 5: Visualization of the six objective and three perceptual dimensions of the MUSHRA-DG test. The objective scores are represented using stacked bars, where multiple error categories are displayed cumulatively rather than as independent percentages. The subjective dimensions are represented using a scatter plot with scores ranging from 0 to 100.
@@ 6.3 Evaluations using MUSHRA-DG-NMR
7 Generalization to English
Thus far, our study has primarily focused on Hindi and Tamil, two low-resource languages from distinct language families, Indo-Aryan and Dravidian, spoken by a combined 562 million people. This choice ensures coverage, diversity, and generalizability across language families in India. We do realize that evaluating our proposed MUSHRA variants on a high-resource language like English would further assess the broader applicability of our findings. Expanding the analysis requires additional resources, as human evaluations are both time and cost intensive. Given constraints, we focus our exploration on MUSHRA and the combined variant, MUSHRA-DG-NMR, drawing parallels with CMOS to analyze how our refinements translate to English. We conducted these evaluations on English TTS systems trained on the LJSpeech dataset, with 30 native US English speakers rating 30 utterances each, balanced across age (18 to 60) and gender. Additionally, we performed a CMOS test, where 15 participants compared all systems against the reference, rating 30 utterances each. Table 7 presents the results, which closely align with our earlier findings in Hindi and Tamil, reinforcing the robustness of our proposed evaluation methodologies.
7.1 Does the Reference-Matching Bias exist?
A key motivation for our work was to assess whether MUSHRA reliably captures the perceptual differences between TTS systems. The scores for English TTS systems in Table 7 reinforce our previous findings: while the reference (REF) is the only system placed in the "Excellent" bin, all other systems are restricted to the "Good" or "Fair" bins. However, CMOS results tell a different story, showing that ST2 and VITS are comparable to, or even preferred over, the reference. This discrepancy highlights MUSHRA’s inability to fully capture perceptual parity or superiority when modern TTS systems approach or exceed reference quality.
These findings further confirm the presence of reference-matching bias (RMB) in English. As a result, MUSHRA scores may misrepresent true perceptual differences, reinforcing the need for evaluation methods that minimize the influence of reference dependence.
7.2 Does Judgement Ambiguity Exist?
To assess the reliability of the mean statistic in MUSHRA for English, we visualize the distribution of MUSHRA scores across raters and systems (see Appendix: Figure 13). These results reaffirm the two key issues previously observed in Tamil and Hindi -
  1. 1.
    High variability in individual ratings: A single rater does not always score a system consistently across different utterances, leading to fluctuations in their assessments.
  2. 2.
    Significant perceptual differences across raters: The mean scores assigned by different raters vary widely, indicating inconsistency in how MUSHRA labels are perceived and applied.
These findings highlight the presence of judgment ambiguity, where raters interpret and use the MUSHRA scale differently, leading to inconsistencies in scoring. This further reinforces our claim that the mean statistic alone is not a reliable measure of perceived quality, as it is strongly influenced by both intra-rater inconsistency and inter-rater variation.
Table 7: Comparison of MUSHRA, MUSHRA-DG-NMR, and CMOS for English.
System MUSHRA MUSHRA-DG-NMR CMOS
µ 95% CI µ 95% CI µ 95% CI
ANC 41.93 3.30 56.38 3.91 - -
FS2 56.86 3.76 63.33 3.32 -0.78 0.22
VITS 72.77 2.75 75.89 2.96 0.02 0.12
ST2 73.27 2.81 78.47 2.84 0.21 0.12
REF 82.62 1.54 79.57 2.96 0 0
7.3 Does our proposed variant (MUSHRA-DG-NMR) help?
Our proposed MUSHRA-DG-NMR test was designed to mitigate reference-matching bias (RMB) and judgment ambiguity, and the English results further confirm its effectiveness. As seen in Table 7, MUSHRA-DG-NMR brings system scores closer to the reference, suggesting that RMB is reduced. At the same time, it preserves the relative rankings of systems while shifting scores, mirroring our findings in Hindi and Tamil. This is particularly important, as it indicates that MUSHRA-DG-NMR does not distort system rankings but instead provides scores that may better reflect perceptual judgments.
While score variance in English does not show a significant reduction, a key strength of MUSHRA-DG-NMR lies in its ability to provide fine-grained insights beyond a single score. This is especially valuable when systems receive similar MUSHRA scores but differ in perceptual quality. By analyzing both subjective and objective dimensions, we can pinpoint the specific factors shaping listener preferences. For example, VITS exhibits nearly twice as many sudden energy variations as ST2, while ST2 has almost double the digital artifacts of VITS. Despite these trade-offs, ST2 is preferred overall due to its improved rhythmic consistency, reinforcing that a simple numerical score does not capture all aspects of TTS quality. For reference, visualizations of both objective and subjective dimensions are provided in Appendix A.7. Although MUSHRA-DG-NMR requires additional time for evaluation, its ability to uncover detailed system strengths and weaknesses makes it a valuable tool for TTS research and development. These fine-grained insights go beyond aggregated scores, offering a deeper understanding of where synthetic speech systems excel and where refinements are needed.
Summary. We release all English ratings, expanding our annotation dataset to 255,000 subjective annotations. These results confirm that MUSHRA scores are affected by reference-matching bias, mean scores can be unreliable due to rating inconsistencies, and MUSHRA-DG-NMR effectively mitigates bias while preserving system rankings. Additionally, our fine-grained analysis provides deeper insights beyond a single score, allowing for a more precise evaluation of system strengths and weaknesses. Overall, these findings further validate our conclusions across both high-resource (English) and low-resource (Hindi and Tamil) languages.
8 Future Work
While our study introduces refined MUSHRA variants to enhance the evaluation of modern TTS systems, several promising directions remain unexplored. The ITU-R standard ITU-R (2015) advises rating no more than 12 signals at a time, yet real-world evaluations could require comparisons across a much larger set of systems for creating a leaderboard. A natural extension of our work would be to adapt MUSHRA-DG-NMR for tournament-style evaluations, allowing new systems to be seamlessly integrated while maintaining a dynamic leaderboard. Another compelling avenue is the development of objective metrics that can either complement subjective evaluations or, in certain cases, serve as viable replacements, offering a more holistic perspective on TTS quality while improving efficiency. Additionally, refining MUSHRA-DG to make it more accessible for large-scale studies could further extend its impact, without sacrificing the fine-grained insights it provides into pronunciation, prosody, and other key speech attributes. Beyond this, exploring cultural and demographic biases within the MUSHRA framework could illuminate potential disparities in how different listener groups perceive synthesized speech, leading to fairer and more representative evaluations. Furthermore, validating our refined MUSHRA variants in practical deployment settings will be crucial to ensure their seamless integration into real-world use cases, fostering broader adoption and iterative refinements based on real-world feedback. These directions build upon our proposed MUSHRA variants and pave the way for more rigorous, equitable, and insightful assessments of TTS systems.
@@ 9 Conclusion
Our comprehensive study reveals significant shortcomings in the current use of the MUSHRA test for evaluating modern high-quality TTS systems. Through an extensive analysis involving 246,000 human ratings for Hindi and Tamil, we identified two primary issues: reference-matching bias and judgment ambiguity. To address these issues, we propose two refined variants of the MUSHRA test: MUSHRA-NMR, which omits explicit identification of the human reference, and MUSHRA-DG, which uses detailed guidelines to calculate MUSHRA scores systematically. Our findings indicate that both variants lead to more reliable evaluations, with MUSHRA-DG offering the additional benefit of fine-grained fault isolation during assessment. Through this work, we also release MANGO, a large human rating dataset, to further support research in this area.
@@ A.1 Instructions For MUSHRA
A.2 CMOS Evidence of Limitations of Reference Matching
The assumption that the human reference always represents the highest quality standard in MUSHRA evaluations may not hold in practice, particularly as modern TTS systems continue to improve. Since MUSHRA inherently relies on reference matching, this approach may unintentionally penalize systems that perform on par with or even surpass the reference. If raters subconsciously anchor their judgments to the provided human reference, systems that differ from it, even in a perceptually superior way, may receive lower scores.
To investigate this, we analyze CMOS-derived preference ratings, where CMOS > 0 indicates a preference for the TTS system over the reference, CMOS < 0 indicates preference for the reference, and CMOS = 0 suggests no preference. As shown in Table 8, StyleTTS2 is preferred over the reference in 54% of cases in Tamil and 43.8% in English, demonstrating that model outputs can surpass human recordings in perceptual quality. Even VITS is preferred 37.2% of times against the reference in English, further challenging the assumption that the human reference should always be treated as the gold standard.
These findings provide direct evidence that MUSHRA’s reliance on reference matching may skew evaluations when systems outperform the reference. This motivates us to explore MUSHRA-NMR, a variant where no explicit reference is provided, allowing for fairer assessments when TTS models reach or exceed human quality. While inspired by multiple listening test methodologies, our goal remains to refine MUSHRA to better accommodate the evolving capabilities of TTS models while preserving its core strengths.
Table 8: CMOS-Derived Preference percentages for TTS Systems in English, Hindi, and Tamil.
Language System Reference (%) Equal (%) System (%)
ST2 39.1 17.1 43.8
VITS 38.7 24.1 37.2
English FS2 68.3 13.5 18.2
ST2 42.8 18.0 39.2
VITS 43.9 18.2 37.8
Hindi FS2 59.9 11.8 28.3
ST2 41.0 5.0 54.0
VITS 63.6 4.8 31.5
Tamil FS2 62.9 3.5 33.5
A.3 Visualizing MUSHRA Distributions
In Section 4, we discussed the distribution of MUSHRA scores across raters for Hindi using Figure 1. Similarly, in Figure 12 and Figure 13, we visualize the MUSHRA scores per rater across the systems for Tamil and English, respectively. Figure 14 and Figure 15, visualizes the MUSHRA scores for each utterance, averaged across raters, for Hindi and Tamil respectively.
A.4 Sensitivity of MUSHRA in Tamil
@@
Figure 9: Summary of participant responses from the post-evaluation survey conducted after the MUSHRA test.
A.5 Post-Evaluation Survey
To better understand how raters made their judgments during the MUSHRA test, we conducted a structured post-evaluation survey with 89 participants after the main evaluation. This survey provided direct insights into participant decision-making, particularly regarding reference-matching bias, judgment ambiguity, rating confusion, and cognitive load. The results, summarized in Figure 9, reinforce key concerns about strict reference dependence in MUSHRA-based evaluations.
A majority of participants (55%) preferred the system that was closer to the reference, confirming that reference-matching plays a strong role in decision-making. However, this bias was not absolute—18% preferred a system that deviated from the reference, and 11% rated solely based on overall quality, demonstrating that strict reference dependence may not always be ideal. Furthermore, responses highlighted judgment ambiguity: while 56% prioritized voice quality over pronunciation errors, 33% favored better pronunciation over voice quality, showing that raters face trade-offs that are not always resolved by aligning with the reference.
Despite the provided instructions, 56% of participants reported experiencing rating confusion at least a few times, confirming that label ambiguity exists in the MUSHRA framework. Notably, 58% of raters did not rate the reference as "Excellent," with 19% rating it as Fair and 1% as Poor, further challenging the assumption that the reference always represents the highest quality standard. Additionally, 22% of participants found the MUSHRA test difficult or very difficult, suggesting that cognitive load may impact rating consistency.
These findings support our argument that while MUSHRA remains a valuable evaluation tool, strict reference-matching and scoring assumptions may not always be ideal. The observed ambiguity, cognitive challenges, and frequent encounters with confusing rating situations, reinforce the need for alternative evaluation setups, such as MUSHRA-NMR, which relaxes reference dependence while maintaining the benefits of comparative evaluation. The survey results directly informed our analysis, ensuring that our interpretations were grounded in participant responses rather than speculation. For additional details, the full survey findings are referenced in Figure 9.
A.6 Detailed Guidelines for MUSHRA-DG
@@
A.7 Fault Isolation in English and Tamil
Similar to Figure 5, we visualize the granular ratings across the objective and perceptual dimensions for English in Figure 11 and, Tamil in Figure 10.
@@ A.7 Fault Isolation in English and Tamil
(a) Objective (English)
(b) Perceptual (English)
A.8 Limitations
@@
Our study focuses on human evaluations for Hindi and Tamil, representing a major Indo-Aryan and Dravidian language, respectively. However, we did not extend our analysis to English, a widely spoken and diverse language. This limitation is due to the scope of our current research and resource constraints. We also wish to state that while the demographics table summarizes the age and gender of participants, details for 21 out of the 492 participants are missing as raters opted out from providing this information. This constitutes a minor fraction and we believe does not significantly affect the analysis. While this work focuses on identifying two key issues — reference-matching bias and judgment ambiguity — several other biases may also exist in TTS evaluations, which warrant further investigation. A more broader analysis could provide more comprehensive insights into the applicability and robustness of our proposed MUSHRA variants across diverse linguistic contexts and practical applications.
A.9 Ethical Considerations
@@
We prioritized ethical conduct throughout our research. All human listeners involved in the study provided informed consent before participating in the evaluation, recruited through professional data annotation agencies. These agencies verified participant language proficiency for task relevance. We established an education criterion of completing high school to ensure participants’ ability to accurately annotate audio content. Participants were compensated fairly for their time and expertise, following industry standards. They were also fully informed about the study nature, procedures, and their right to withdraw at any point without consequence. We ensured that our study adhered to ethical guidelines by obtaining approval from the Institutional Ethics Committee, which reviewed our methodology and confirmed compliance with ethical standards.
@@
We strived for inclusivity and bias mitigation. Participants came from diverse demographic backgrounds, and we recruited only native speakers to capture the subtle linguistic and cultural nuances of each language. To minimize rater burden and bias in the new MUSHRA test variations, we prioritized user-friendliness and transparency in the design, providing clear guidelines.
@@
We release the evaluations dataset under CC-BY-4.0 license after careful consideration of privacy and ethical use. Identifiable information about the participants was anonymized to protect their privacy. We encourage the use of this dataset for advancing TTS evaluation metrics, emphasizing that it should be used responsibly and ethically, adhering to principles of transparency and fairness. Finally, we acknowledge that our study focuses on Hindi and Tamil, and we recognize the importance of extending such evaluations to other languages, including English, to generalize our findings. Future research should continue to explore these ethical dimensions, ensuring that the development and evaluation of TTS systems are conducted with respect for the diversity and rights of all participants involved.
@@
Figure 13: Visualization of the MUSHRA score distributions per rater across three systems— FS2, ST2, and VITS, along with the reference (REF) and anchor (ANC) for English. Each boxplot represents ratings (0-100) across all test utterances for a system by one rater. The substantial heights of some boxplots indicate significant variance in the scores of that rater. The variation in boxplot means across raters suggests a high level of inter-rater variance